# Synthetic Data for Robust Stroke Segmentation

### v2

@@ Appendix A Further segmentation metrics
T1w Baseline 0.646 24.5 18.04 6.19 0.335 0.600 1.178
Baseline+TTA 0.648 23.7 18.33 3.42 0.442 0.598 1.126
Baseline+TENT 0.077 144.9 75.50 3.51 0.180 0.060 0.285
Baseline+DAE 0.614 26.2 26.53 3.15 0.507 0.685 2.238
Baseline+PL 0.325 67.5 146.72 175.61 0.017 0.613 10.342
Baseline+UPL 0.341 74.4 48.10 22.84 0.244 0.262 0.695
Baseline+DPL 0.340 52.5 52.48 9.53 0.324 0.247 0.465
Synth 0.609 27.9 22.88 1.99 0.613 0.539 0.953
Synth+TTA 0.611 27.0 24.08 1.84 0.640 0.534 0.885
Synth+TENT 0.560 33.4 26.71 2.42 0.528 0.604 2.385
Synth+DAE 0.613 23.4 20.60 2.26 0.581 0.640 1.903
Synth+PL 0.311 158.6 48.41 260.54 0.017 0.310 2.303
Synth+UPL 0.340 56.3 54.57 20.35 0.334 0.243 0.457
Synth+DPL 0.255 58.3 63.36 7.44 0.530 0.169 0.339
T2w Baseline 0.040 76.8 73.80 47.12 0.044 0.063 6.778
Baseline+TTA 0.023 77.0 71.83 26.00 0.054 0.036 6.819
Baseline+TENT 0.000 148.5 85.40 5.11 0.047 0.000 0.206
Baseline+DAE 0.056 75.8 123.49 10.09 0.113 0.134 10.938
Baseline+PL 0.006 78.8 127.78 353.31 0.007 0.014 12.731
Baseline+UPL 0.000 256.0 85.74 2.57 0.013 0.000 0.185
Baseline+DPL 0.000 256.0 85.74 2.57 0.013 0.000 0.185
Synth 0.299 66.2 40.79 38.81 0.060 0.304 2.282
Synth+TTA 0.284 66.2 43.70 16.00 0.117 0.274 1.893
Synth+TENT 0.321 55.5 43.43 15.09 0.140 0.326 1.663
Synth+DAE 0.437 40.8 40.19 11.43 0.196 0.410 1.344
Synth+PL 0.384 65.5 40.26 22.77 0.095 0.475 3.881
Synth+UPL 0.120 50.7 73.30 22.71 0.172 0.074 0.484
Synth+DPL 0.055 52.3 78.43 47.28 0.108 0.032 0.447
FLAIR Baseline 0.193 66.2 52.15 49.49 0.049 0.205 4.532
Baseline+TTA 0.202 65.7 47.98 25.98 0.090 0.197 3.719
Baseline+TENT 0.011 165.6 79.40 4.94 0.070 0.008 0.628
Baseline+DAE 0.287 54.8 75.47 11.09 0.232 0.327 5.981
Baseline+PL 0.126 81.9 162.13 382.66 0.007 0.240 13.548
Baseline+UPL 0.000 256.0 82.36 2.62 0.035 0.000 0.497
Baseline+DPL 0.000 256.0 82.36 2.62 0.035 0.000 0.497
Synth 0.199 59.3 60.36 17.69 0.150 0.157 1.349
Synth+TTA 0.190 56.4 62.98 8.36 0.256 0.143 1.185
Synth+TENT 0.340 49.7 46.05 5.01 0.247 0.286 1.582
Synth+DAE 0.263 51.5 53.14 5.88 0.248 0.215 1.581
Synth+PL 0.276 96.8 45.97 93.18 0.027 0.255 2.133
Synth+UPL 0.029 99.1 80.44 4.79 0.272 0.016 0.520
Synth+DPL 0.010 146.9 81.38 7.11 0.189 0.005 0.512
Ensemble Baseline 0.107 53.3 77.04 22.20 0.309 0.071 1.137
Baseline+TTA 0.087 65.8 78.91 9.10 0.429 0.057 1.126
Baseline+TENT 0.003 238.0 85.73 2.60 0.054 0.002 0.185
Baseline+DAE 0.212 51.3 73.84 39.25 0.168 0.212 2.350
Baseline+PL 0.016 76.1 70.76 512.92 0.004 0.014 3.869
Baseline+UPL 0.000 256.0 85.74 2.57 0.013 0.000 0.185
Baseline+DPL 0.000 256.0 85.74 2.57 0.013 0.000 0.185
Synth 0.525 34.9 35.60 10.09 0.356 0.451 0.770
Synth+TTA 0.530 33.0 36.51 4.50 0.470 0.452 0.696
Synth+TENT 0.451 39.1 38.50 11.12 0.370 0.412 0.960
Synth+DAE 0.579 27.6 27.52 6.77 0.432 0.531 0.923
Synth+PL 0.529 50.2 33.21 29.02 0.106 0.515 1.551
Synth+UPL 0.244 53.0 65.77 9.80 0.470 0.159 0.310
Synth+DPL 0.172 71.4 71.62 13.97 0.414 0.107 0.284
Table 14: Mean results on the PLORAS dataset (N=263). Best score per column is shown in bold. FPR values are shown as a multiple of 1000.
Modality Model Dice HD95 AVD ALD LF1 TPR FPR
T2 Baseline 0.017 79.8 116.97 11.99 0.108 0.046 8.483
Baseline+TTA 0.011 78.8 94.22 11.14 0.101 0.034 7.071
Baseline+TENT 0.001 98.7 21.88 5.20 0.019 0.006 0.478
Baseline+DAE 0.026 83.6 206.65 3.28 0.269 0.097 13.761
Baseline+PL 0.004 76.8 51.10 19.19 0.032 0.009 3.874
Baseline+UPL 0.000 256.0 22.47 3.29 0.000 0.005 0.414
Baseline+DPL 0.000 256.0 22.47 3.29 0.000 0.005 0.414
Synth 0.188 74.5 43.14 11.42 0.227 0.323 3.083
Synth+TTA 0.195 72.3 38.65 6.56 0.310 0.314 2.771
Synth+TENT 0.153 74.7 70.50 19.57 0.119 0.265 4.774
Synth+DAE 0.208 63.3 29.07 5.42 0.280 0.275 2.008
Synth+PL 0.077 65.3 23.96 3.78 0.232 0.079 1.300
Synth+UPL 0.014 168.9 21.76 2.75 0.124 0.012 0.420
Synth+DPL 0.027 110.9 20.14 2.94 0.097 0.021 0.489
FLAIR Baseline 0.005 77.9 37.54 19.76 0.028 0.010 2.930
Baseline+TTA 0.002 78.1 30.08 12.72 0.020 0.006 2.169
Baseline+TENT 0.002 105.1 21.69 5.15 0.010 0.003 0.647
Baseline+DAE 0.014 78.6 103.77 4.12 0.145 0.043 7.639
Baseline+PL 0.003 81.7 102.92 118.06 0.015 0.011 7.674
Baseline+UPL 0.000 256.0 22.72 3.05 0.000 0.002 0.537
Baseline+DPL 0.000 256.0 22.72 3.05 0.000 0.002 0.537
Synth 0.213 75.8 48.25 17.65 0.131 0.350 3.374
Synth+TTA 0.228 74.7 40.69 8.99 0.189 0.341 2.874
Synth+TENT 0.149 85.8 44.09 9.20 0.111 0.197 3.080
Synth+DAE 0.210 68.7 41.66 8.85 0.170 0.268 3.082
Synth+PL 0.190 90.7 92.65 23.32 0.131 0.437 6.027
Synth+UPL 0.223 77.3 21.27 6.00 0.231 0.253 1.747
Synth+DPL 0.048 111.4 21.41 3.05 0.207 0.031 0.584

### v3

@@ Appendix B Further segmentation metrics
Modality Model Dice HD95 AVD ALD LF1 TPR FPR
T1w Baseline 0.646 24.5 18.04 6.19 0.335 0.600 1.178
Baseline+TTA 0.648 23.7 18.33 3.42 0.442 0.598 1.126
Baseline+TENT 0.077 144.9 75.50 3.51 0.180 0.060 0.285
Baseline+DAE 0.614 26.2 26.53 3.15 0.507 0.685 2.238
Baseline+PL 0.325 67.5 146.72 175.61 0.017 0.613 10.342
Baseline+UPL 0.341 74.4 48.10 22.84 0.244 0.262 0.695
Baseline+DPL 0.340 52.5 52.48 9.53 0.324 0.247 0.465
Synth (Ours) 0.609 27.9 22.88 1.99 0.613 0.539 0.953
Synth+TTA 0.611 27.0 24.08 1.84 0.640 0.534 0.885
Synth+TENT 0.560 33.4 26.71 2.42 0.528 0.604 2.385
Synth+DAE 0.613 23.4 20.60 2.26 0.581 0.640 1.903
Synth+PL 0.311 158.6 48.41 260.54 0.017 0.310 2.303
Synth+UPL 0.340 56.3 54.57 20.35 0.334 0.243 0.457
Synth+DPL 0.255 58.3 63.36 7.44 0.530 0.169 0.339
T2w Baseline 0.040 76.8 73.80 47.12 0.044 0.063 6.778
Baseline+TTA 0.023 77.0 71.83 26.00 0.054 0.036 6.819
Baseline+TENT 0.000 148.5 85.40 5.11 0.047 0.000 0.206
Baseline+DAE 0.056 75.8 123.49 10.09 0.113 0.134 10.938
Baseline+PL 0.006 78.8 127.78 353.31 0.007 0.014 12.731
Baseline+UPL 0.000 256.0 85.74 2.57 0.013 0.000 0.185
Baseline+DPL 0.000 256.0 85.74 2.57 0.013 0.000 0.185
Synth (Ours) 0.299 66.2 40.79 38.81 0.060 0.304 2.282
Synth+TTA 0.284 66.2 43.70 16.00 0.117 0.274 1.893
Synth+TENT 0.321 55.5 43.43 15.09 0.140 0.326 1.663
Synth+DAE 0.437 40.8 40.19 11.43 0.196 0.410 1.344
Synth+PL 0.384 65.5 40.26 22.77 0.095 0.475 3.881
Synth+UPL 0.120 50.7 73.30 22.71 0.172 0.074 0.484
Synth+DPL 0.055 52.3 78.43 47.28 0.108 0.032 0.447
FLAIR Baseline 0.193 66.2 52.15 49.49 0.049 0.205 4.532
Baseline+TTA 0.202 65.7 47.98 25.98 0.090 0.197 3.719
Baseline+TENT 0.011 165.6 79.40 4.94 0.070 0.008 0.628
Baseline+DAE 0.287 54.8 75.47 11.09 0.232 0.327 5.981
Baseline+PL 0.126 81.9 162.13 382.66 0.007 0.240 13.548
Baseline+UPL 0.000 256.0 82.36 2.62 0.035 0.000 0.497
Baseline+DPL 0.000 256.0 82.36 2.62 0.035 0.000 0.497
Synth (Ours) 0.199 59.3 60.36 17.69 0.150 0.157 1.349
Synth+TTA 0.190 56.4 62.98 8.36 0.256 0.143 1.185
Synth+TENT 0.340 49.7 46.05 5.01 0.247 0.286 1.582
Synth+DAE 0.263 51.5 53.14 5.88 0.248 0.215 1.581
Synth+PL 0.276 96.8 45.97 93.18 0.027 0.255 2.133
Synth+UPL 0.029 99.1 80.44 4.79 0.272 0.016 0.520
Synth+DPL 0.010 146.9 81.38 7.11 0.189 0.005 0.512
Ensemble Baseline 0.107 53.3 77.04 22.20 0.309 0.071 1.137
Baseline+TTA 0.087 65.8 78.91 9.10 0.429 0.057 1.126
Baseline+TENT 0.003 238.0 85.73 2.60 0.054 0.002 0.185
Baseline+DAE 0.212 51.3 73.84 39.25 0.168 0.212 2.350
Baseline+PL 0.016 76.1 70.76 512.92 0.004 0.014 3.869
Baseline+UPL 0.000 256.0 85.74 2.57 0.013 0.000 0.185
Baseline+DPL 0.000 256.0 85.74 2.57 0.013 0.000 0.185
Synth (Ours) 0.525 34.9 35.60 10.09 0.356 0.451 0.770
Synth+TTA 0.530 33.0 36.51 4.50 0.470 0.452 0.696
Synth+TENT 0.451 39.1 38.50 11.12 0.370 0.412 0.960
Synth+DAE 0.579 27.6 27.52 6.77 0.432 0.531 0.923
Synth+PL 0.529 50.2 33.21 29.02 0.106 0.515 1.551
Synth+UPL 0.244 53.0 65.77 9.80 0.470 0.159 0.310
Synth+DPL 0.172 71.4 71.62 13.97 0.414 0.107 0.284
@@
Table 13: Mean results on the PLORAS dataset (N=661). Best score per column is shown in bold. FPR values are shown as a multiple of 1000.
@@
T2w Baseline 0.028 80.3 112.82 12.22 0.101 0.064 9.638
Baseline+TTA 0.020 79.1 92.69 11.02 0.103 0.045 8.409
Baseline+TENT 0.005 101.5 30.79 5.12 0.020 0.007 1.849
Baseline+DAE 0.044 82.7 204.49 2.64 0.276 0.123 14.983
Baseline+PL 0.000 235.8 32.46 2.13 0.009 0.004 1.696
Baseline+UPL 0.000 256.0 32.47 2.12 0.009 0.004 1.695
Baseline+DPL 0.000 256.0 32.47 2.12 0.009 0.004 1.695
Synth (Ours) 0.250 71.9 40.20 8.42 0.243 0.371 4.156
Synth+TTA 0.253 70.5 37.00 4.75 0.329 0.361 3.917
Synth+TENT 0.202 77.0 61.99 21.23 0.118 0.330 5.612
Synth+DAE 0.291 58.7 23.68 4.28 0.307 0.312 2.969
Synth+PL 0.162 81.0 33.28 8.58 0.179 0.186 3.737
Synth+UPL 0.101 66.0 25.80 11.07 0.186 0.073 2.014
Synth+DPL 0.053 72.1 28.44 3.42 0.212 0.036 1.832
FLAIR Baseline 0.010 74.2 42.12 18.69 0.033 0.013 2.839
Baseline+TTA 0.006 74.6 38.04 12.57 0.026 0.008 2.119
Baseline+TENT 0.000 106.3 37.65 4.99 0.014 0.003 0.718
Baseline+DAE 0.027 76.2 98.18 3.81 0.155 0.055 7.805
Baseline+PL 0.000 249.9 39.25 2.46 0.000 0.002 0.582
Baseline+UPL 0.000 256.0 39.25 2.53 0.000 0.002 0.582
Baseline+DPL 0.000 256.0 39.25 2.53 0.000 0.002 0.582
Synth (Ours) 0.296 72.1 48.04 17.93 0.137 0.415 3.443
Synth+TTA 0.314 71.4 41.15 9.07 0.213 0.412 2.967
Synth+TENT 0.221 79.4 48.21 9.87 0.140 0.280 3.362
Synth+DAE 0.304 63.6 43.34 8.82 0.199 0.359 3.258
Synth+PL 0.279 84.3 81.32 20.01 0.156 0.519 5.573
Synth+UPL 0.145 78.5 31.51 15.50 0.115 0.108 1.028
Synth+DPL 0.243 80.1 25.71 13.82 0.146 0.224 1.514
CT Baseline 0.029 70.9 28.01 8.54 0.084 0.032 1.492
Baseline+TTA 0.017 125.0 27.63 4.11 0.076 0.011 0.593
Baseline+TENT 0.000 207.7 28.74 2.59 0.018 0.000 0.550
Baseline+DAE 0.064 78.2 78.65 2.98 0.198 0.106 5.426
Baseline+PL 0.011 78.3 79.93 2010.46 0.002 0.032 6.333
Baseline+UPL 0.000 256.0 28.87 2.84 0.000 0.000 0.542
Baseline+DPL 0.000 256.0 28.87 2.84 0.000 0.000 0.542
Synth (Ours) 0.234 66.8 19.71 5.43 0.233 0.204 0.965
Synth+TTA 0.229 85.2 20.34 2.65 0.317 0.195 0.913
Synth+TENT 0.222 58.6 20.82 9.17 0.139 0.210 1.245
Synth+DAE 0.307 51.6 23.81 4.13 0.258 0.298 1.657
Synth+PL 0.000 110.0 26.69 34.56 0.002 0.001 0.714
Synth+UPL 0.000 255.0 28.86 2.84 0.000 0.000 0.543
Synth+DPL 0.000 256.0 28.87 2.84 0.000 0.000 0.542
Appendix C Wilcoxon Significance Tests
Significance measures are provided below in Figures 7 - 17 as Wilcoxon signed-rank test values for pairwise comparisons between models on each dataset. ’Oracle DA’ represents the hypothetical best-case scenario where optimal DA method is known a priori for each dataset/modality and applied to the baseline model. Median Dice values are shown along the diagonal.
Figure 7: Wilcoxon signed-rank test values for Dice metric measurements in the ATLAS T1w dataset.
Figure 8: Wilcoxon signed-rank test values for Dice metric measurements in the ARC T1w dataset.
Figure 9: Wilcoxon signed-rank test values for Dice metric measurements in the ARC T2w dataset.
Figure 10: Wilcoxon signed-rank test values for Dice metric measurements in the ARC FLAIR dataset.
Figure 11: Wilcoxon signed-rank test values for Dice metric measurements in the ISLES 2015 T1w dataset.
Figure 12: Wilcoxon signed-rank test values for Dice metric measurements in the ISLES 2015 T2w dataset.
Figure 13: Wilcoxon signed-rank test values for Dice metric measurements in the ISLES 2015 FLAIR dataset.
Figure 14: Wilcoxon signed-rank test values for Dice metric measurements in the ISLES 2015 DWI dataset.
Figure 15: Wilcoxon signed-rank test values for Dice metric measurements in the PLORAS T2w dataset.
Figure 16: Wilcoxon signed-rank test values for Dice metric measurements in the PLORAS FLAIR dataset.
Figure 17: Wilcoxon signed-rank test values for Dice metric measurements in the PLORAS CT dataset.
Appendix D Additional Qualitative Results
Results are shown below in Figures 18 - 23 for failure and success cases for the three different sequences in the PLORAS dataset.
Figure 18: Sample visualisations of successful cases in the PLORAS T2w dataset. Green indicates a true positive prediction, red a false positive, and blue a false negative.
Figure 19: Sample visualisations of failure cases in the PLORAS T2w dataset. Green indicates a true positive prediction, red a false positive, and blue a false negative.
Figure 20: Sample visualisations of successful cases in the PLORAS FLAIR dataset. Green indicates a true positive prediction, red a false positive, and blue a false negative.
Figure 21: Sample visualisations of failure cases in the PLORAS FLAIR dataset. Green indicates a true positive prediction, red a false positive, and blue a false negative.
Figure 22: Sample visualisations of successful cases in the PLORAS CT dataset. Green indicates a true positive prediction, red a false positive, and blue a false negative.
Figure 23: Sample visualisations of failure cases in the PLORAS CT dataset. Green indicates a true positive prediction, red a false positive, and blue a false negative.